OpenStudy (anonymous):

A uniform disk with mass 41.7kg and radius 0.290m is pivoted at its center about a horizontal, frictionless axle that is stationary. The disk is initially at rest, and then a constant force 28.0N is applied tangent to the rim of the disk. What is the magnitude v of the tangential velocity of a point on the rim of the disk after the disk has turned through 0.120 revolution? What is the magnitude a of the resultant acceleration of a point on the rim of the disk after the disk has turned through 0.120 revolution?

OpenStudy (anonymous):

Strategy: 1) Find angular acceleration alpha first. 2) Find angular velocity omega after 0.120 rev 3) Find tangential velocity from omega found in (2) 4)

OpenStudy (anonymous):

oh okay ill equate it

OpenStudy (anonymous):

@michelle092294 : you find the answer?

OpenStudy (anonymous):

no -_- i keep having the wrong answer. then i got 1 more chance. and ill have zero. -_- can you send me the equations?

OpenStudy (anonymous):

tangential velocity = radius * omega

OpenStudy (anonymous):

What is the moment oh inertia of disk with mass m radius r?

OpenStudy (anonymous):

\[V=r* \omega\] where omega= angular velocity

OpenStudy (anonymous):

now find omega first

OpenStudy (anonymous):

thank you!!!! :D

OpenStudy (anonymous):

\[torque= inertia* \omega \]

OpenStudy (anonymous):

inertia= m r^2

OpenStudy (anonymous):

thank you jason :)

OpenStudy (anonymous):

:) You're Welcome michi

OpenStudy (anonymous):

\[(\omega _{f})^2=2*\alpha*\theta\] initial angular velocity = 0 okay?

OpenStudy (anonymous):

okay :D now il solve it :)

OpenStudy (anonymous):

great

OpenStudy (anonymous):

\[\alpha= equation ?\]

OpenStudy (anonymous):

\[\tau=I*\alpha=F*radius\] where I is moment of inertia and F= force

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

michi , find inertia , you have force and radius ..you'll get alpha ..after that you can calculate angular velocity ...

OpenStudy (anonymous):

yes got the inertia value. now need to get alpha :)